Fill in the book value at the end of year 1 under each depreciation method when residual value is $36,000 and useful life is 4 years:

Depreciation method:

Straight-line

$335,250

Units-of-Output

Double-declining

$217,500
Answer #1

Using Straight Line method:

Book Value= $335,250

Residual Value= $36,000

Value to be depreciated = 335,250- 36,000= 299,250

Life= 4 years

depreciation each year= 299250/4=

$74,812.5

book value at the end of year1= 335,250-74,812.5

= $260,437.5

Double declining balance method:

Book Value= $217,500

Straigth line dep percent= 100%/4= 25%

Depreciation = 2* Straigth line dep percent * Book Value

=2*0.25*217,500

= 108,750.

Book Value at the end of year 1= 217500-108750

=108,750